When 0.070 mol of propionic acid, C2H5CO2H, is dissolved in 450 mL of water, the equilibrium concentration of H3O+ ions is measured to be 2.19 x 10−3 M. What is Ka for this acid?
Answer #1

[C2H5CO2H] = mol of C2H5CO2H / volume in L
= 0.070 mol / 0.450 L
= 0.156 M

C2H5CO2H dissociates as:

C2H5CO2H          ----->     H+   + C2H5CO2-
0.156                 0         0
0.156-x               x         x


Ka = [H+][C2H5CO2-]/[C2H5CO2H]
Ka = x*x/(c-x)
Ka = 2.19*10^-3*2.19*10^-3/(0.156-2.19*10^-3)
Ka = 3.118*10^-5
Answer: 3.12*10^-5